# A Detection of Circumgalactic Dust at Megaparsec Scales with Maximum Likelihood Estimation

### v1

@@ 3.2 Star catalogs for systematics testing
To create a background star catalog (for cross-correlation with the WISExSuperCOSMOS foreground), we query the DES Y3 GOLD photometric catalog using the star selection criteria in I. Sevilla-Noarbe et al. (2021). An additional cut of $18<$ mag_r_corr $<21$ eliminates saturated stars and possible galaxy confusion. The final catalog sizes are 824,345 stars in the “signal” catalog and 9,892,452 stars in the “random” catalog. Because these catalogs contain the same magnitude measures as the joined redMaGiC-Y3 GOLD galaxy, comparison of results between the two is very straightforward.
4 Methods
This section presents the analysis pipeline we put in place to apply the formalism of Section 2 and so measure circumgalactic dust extinction profiles. The major steps of the pipeline are: intake and processing of foreground and background catalogs including random catalogs; defining a particular dust extinction model; dividing of galaxies into redshift bins and de-meaning of galaxy colors; per-galaxy calculation of the value and uncertainty of $A_{V}$; calculation of $\xi_{NA_{V}}(\theta)$; and finally, the fitting of a polynomial to the resulting extinction profile. Details are provided in the subsections below. The code developed to carry out this analysis, dusthalos, is hosted on GitHub [^4] https://github.com/mcclearyj/dusthalos/tree/master and will be made publicly available upon acceptance of this manuscript.
4.1 Catalog handling
@@
The specific implementation of this workflow for the galaxy and star catalogs used in our analysis is mostly covered in Section 3, with a few additional details below.
Masking. To ensure an exact match in sky coverage of the foreground and background catalogs, we use healpy software (A. Zonca et al., 2019), a Python implementation of the Hierarchical Equal Area isoLatitude Pixelization (HEALPix [^5] http://healpix.sourceforge.net, K. M. Górski et al., 2005) scheme to manage discrete data on the sphere. Although HEALPix survey masks are available for the DES Y3 and WISExSuperCOSMOS catalogs, they are not available for Gaia. Accordingly, we have added to our dust_halos code the ability to make a survey mask of a desired $N_{\mathrm{SIDE}}$ parameter (resolution) given a set of cataloged celestial coordinates. First, $\alpha,\delta$ are converted to colatitude and longitude coordinates, then the healpy ang2pix method is used to assign these to HEALPix pixels on a map of resolution set by the given NSIDE parameter. A Python NumPy array is initialized with the number of elements set with the HEALPy nside2npix method. Empty HEALPix pixels are set to zero, HEALPix pixels that include stars/galaxies are set to 1. The array is then saved to file and manipulated as normal.
@@
Photometry. The redMaGiC catalogs we use in our analysis include celestial coordinates and redshifts but not magnitudes. As described in Section 3, to obtain colors for the real redMaGiC catalogs, we join member galaxies to their counterparts in the Y3 GOLD catalog on the COADD_OBJECT_ID parameter. Random catalog entries lack a corresponding entry in the Y3 GOLD catalog and so are randomly assigned photometry as follows:
  1. 1.
    Bin the random catalog and reference photometric catalog by redshift.
  2. 2.
    For each redshift bin, generate an array of random integers with length equal to the number of random galaxies and with an upper range set to the number of photometric-catalog galaxies.
  3. 3.
    Assign the $g$, $r$, $i$, $z$ magnitudes of galaxies in the photometry catalog to the corresponding entries in the random catalog, element-wise.
The color-matched random galaxy catalogs resulting from this procedure are shown in Figure 2. Aside from a small offset in the last bin of the $g-r$ and $i-z$ color space for the high-redshift sample, the real and random galaxy catalogs are closely matched. Note that the photometric catalogs used for this procedure are the color-augmented redMaGiC galaxy catalogs, not the full Y3 GOLD catalog, as redMaGiC galaxies occupy a unique color space. That said, we found that adopting color-redshift-matched randoms versus randomly assigned colors from the full GOLD catalog did not noticeably affect extinction profiles.
@@
4.2 Extinction calculation and correlation
To obtain values for $d\mathrm{M}/d\rho$, the model for dust-grain induced reddening referenced in Equations 35, we use the comprehensive dust extinction law published by K. D. Gordon et al. (2023), hereafter G23. The G23 model requires specification of the wavelengths at which to compute the extinction and a value for $R_{V}$. For the former, we use the effective wavelengths of the DECam filter set available through the Virtual Observatory’s Filter Profile Service, rescaled as prescribed in Table 6 of E. F. Schlafly & D. P. Finkbeiner (2011): $\left\{g,r,i,z\right\}=\left\{4796.6,6382.6,7769.0,9108.2\right\}$. For the latter, in the absence of a compelling reason to do otherwise, we adopt the common default of $R_{V}=3.1$. (Further consideration is given to the choice of $R_{V}$ in Section 6.)
@@ 4.2 Extinction calculation and correlation
We compute $\xi_{NA_{V}}(\theta)$ using the galaxy-kappa cross-correlation method NKCorrelation provided in the TreeCorr package [^6] https://github.com/rmjarvis/TreeCorr (M. Jarvis et al., 2004). Though intended to correlate galaxy counts with convergence (i.e., $\kappa$) the code is flexible enough to accommodate any scalar quantity such as $A_{V}$. All correlations are taken in 13 impact parameter bins, from 0.5 to 200 arcminutes, or about 20 kiloparsecs to 10 megaparsecs in the average foreground galaxy rest frame. This range matches the physical scales probed in M10 and related works. Error bars on extinction profiles are based on TreeCorr’s sample covariance estimation method, which is the simplest method available in that package to account for the combination of Poisson counting statistics, statistical variance of the scalar quantity, and sample variance across the observation.
The final step of our pipeline is the fitting of power laws of the form $A_{V}(r)=Cr^{\alpha}$ to the extinction profiles. Although halos are more complex than this simple parametrization suggests, it is nevertheless a convenient heuristic for comparison across galaxy samples and with the literature. We perform weighted least-squares fits to the extinction profiles with the Python package StatsModel.
Our pipeline enables several tests for systematic error. Substitution of star catalogs for galaxy foregrounds and backgrounds offers a valuable null test for our pipeline, as main-sequence stars do not have extended halos of dust. Similarly, stars embedded in the Milky Way should not display dust-induced color gradients as a function of projected distance from galaxies outside the Milky Way. Additionally, inverting foreground and background samples, i.e., using a high-redshift catalog as a foreground and a low-redshift catalog as a background, should result in a non-detection of reddening. We deploy all three of these tests in the course of our analysis.
5 Results
@@
$A^{\mathrm{M10}}_{V}(r)=\left(4.4\pm 1.1\right)\times 10^{-3}\left(\frac{r}{100h^{-1}\,\mathrm{kpc}}\right)^{-0.86\pm 0.19}$ (9)
@@
5.1 Galaxy extinction profiles
@@
The extinction profiles of WISExSuperCOSMOS galaxies are shown in Figure 3: the result from cross-correlation with the $z>0.5$ high-density redMaGiC sample is plotted in dark blue, and the result from cross-correlation with the high-redshift redMaGiC sample is plotted in light blue. Both extinction profiles follow a broken power law, with a rapid decrease from $A_{V}\sim 0.3$ to $A_{V}\sim 0.02$ in the inner 50 $h^{-1}\,$kpc, then a slower decline to $A_{V}\sim 0.02$ at 1 $h^{-1}\,$pc. At the transition to the 2-halo regime ($r\gtrsim 2$ $h^{-1}\,$pc), the extinction profiles rapidly diminish to below our detection threshold ($A_{V}\sim 10^{-5}$).
@@
We fit a power law to the WISExSuperCOSMOS extinction profiles in the 1-halo regime (50 $h^{-1}\,$kpc $<r<1$ $h^{-1}\,$pc), where the lower limit is set to the point where the redMaGiC LRG and WISExSuperCOSMOS extinction profiles intersect, and the upper limit is set to the point where the extinction profiles taper off. Although the selection functions of the two redMaGiC background catalogs are different, the foreground catalog is of course the same. We thus fit a single power law to the two WISExSuperCOSMOS $\times$ redMaGiC profiles simultaneously, producing Equation 10. Quoted uncertainties reflect the standard error on the mean parameter values.
$A_{V,\,\mathrm{stacked}}(r)=\left(3.1\pm 0.06\right)\times 10^{-3}\left(\frac{r}{100\,h^{-1}\,{\rm kpc}}\right)^{-0.87\pm 0.04}$ (10)
@@
Cross-correlation of the redMaGiC LRG galaxy sample with the redMaGiC high-luminosity high-$z$ sample yields the extinction profile plotted in maroon in Figure 3. As might be expected for this galaxy population, there is little apparent dust at the lowest impact parameters, within a few optical radii of the galaxies ($r<50$ $h^{-1}\,$kpc). Between $20.05<r<$ 2 $h^{-1}\,$pc, the LRG profile is qualitatively similar to the WISExSCOS profiles. However, the LRG sample shows significant extinction out to 30 Mpc, with a distinct dip at $\sim 2$ $h^{-1}\,$pc—the point at which the WISExSCOS profiles taper off. We fit a power law to the LRG $A_{V}$ profile in the range 50 $h^{-1}\,$kpc $<r<30$ $h^{-1}\,$pc, with the result given in Equation 11. As before, quoted uncertainties reflect the standard error on the mean parameter values.
$A_{V,\,\mathrm{LRG}}(r)=\left(4.6\pm 0.2\right)\times 10^{-3}\left(\frac{r}{100\,h^{-1}\,\mathrm{kpc}}\right)^{-0.77\pm 0.04}$ (11)
5.2 Systematics tests
@@
Figure 4 presents the results of the systematics tests described in Section 4, namely, the calculation of extinction profiles using star catalogs and with inverted foreground and background samples (high-$z$ foreground, low-$z$ background). The top row of Figure 4 shows the the extinction profile resulting from the cross-correlation of the Gaia star foreground catalog with the redMaGiC high-luminosity high-redshift sample (maroon points) and high-density sample (red points). The middle row of Figure 4 shows the $A_{V}$ profile resulting from the cross-correlation of the WISExSuperCOSMOS foreground with the background catalog created from DES Y3 stars. For the reverse $A_{V}$ profile test, we invert the redMaGiC LRG measurement and cross-correlate the high-redshift redMaGiC catalog with the $z<0.45$ high-density redMaGiC catalog. The outcome is shown in the bottom row of Figure 4.
@@
As with the galaxy samples, we fit power laws to the systematics-test extinction profiles, including the combined Gaia-redMaGiC high-density and high-redshift profiles. The resulting extinction profile parameters are summarized in Table 1 and also plotted in Figure 4. Weighted least squares fits to these extinction profiles are effectively all consistent with zero, with prefactors close to unity and exponents on the order of $10^{-5}$. The only fit that yielded a moderately significant ($3\,\sigma$) exponent was for the inverted foreground-background $A_{V}$ profile. Even so, the magnitude of this exponent is only one ten-thousandth of the magnitude of the galaxy $A_{V}$ profile exponents, indicating no cause for concern.
Table 1: Systematics test $A_{V}$ profile fits
Sample Prefactor Exponent (×$\times 10^{-5}$)
Gaia $\times$ redMaGiC high-density $1.0\,\pm\,(8.0\times 10^{-5})$ $-1.0\,\pm$ 1.7
[…]
Gaia $\times$ redMaGiC combined $1.0\,\pm\,(6.0\times 10^{-5}$) $-0.7\pm 1.3$
[…]
@@ 6 Discussion
Figure 3 shows an unambiguous detection of $A_{V}$ at megaparsec scales about both redMaGiC LRGs and a mixed sample of quiescent and star-forming galaxies in WISExSCOS, with negligible contributions from systematic errors (Figure 4 and Table 1). The slopes ($\alpha\simeq{0.8}$) of the extinction profiles presented in Equations 10 and 11 are in excellent agreement with M10 (cf. Equation 9) and other studies both EW(Mg II)-based (G. Zhu et al., 2014; T.-W. Lan & H. Mo, 2018; Z. Chen et al., 2024) and extinction-based (Z. Ruoyi & Y. Haibo, 2020). This ubiquitous value of $\alpha\simeq 0.8$ is directly connected to halo mass profiles, where the so-called differential surface density typically follows $\Delta\Sigma\propto r^{\beta}$. For massive, dynamically relaxed galaxies, $0.7\leq\beta\leq 0.9$ (F. Brimioulle et al., 2013; M. Velander et al., 2014; E. van Uitert et al., 2016). These values of $\beta$ match both our measured $\alpha\simeq 0.8$ and values of $\alpha$ in the literature, which implies that dust permeates the entire galaxy halo. In addition to the exponent, the prefactor of our $A_{V,\,\mathrm{LRG}}$ is also in excellent agreement with M10, while the $A_{V,\,\mathrm{stacked}}$ is about $1\,\sigma$ lower. Although a direct comparison of our result to Mg II absorbers is impossible without spectroscopic calibration of our samples, by applying the B. Ménard et al. (2008) scaling between $E(B-V)$ and EW(Mg II), we find a (very) approximate EW of 1.4 Å at an impact parameter of 20 kpc for our WISExSCOS sample, within a factor of 2 of the ELG results of T.-W. Lan & H. Mo (2018) and Z. Chen et al. (2024).
@@
We may convert our observed $A_{V}$ profiles into an approximate dust mass surface density profile $\Sigma_{d}(r)$ using the astrodust+PAH model of Milky Way dust described in B. S. Hensley & B. T. Draine (2023). Using the astrodust surface dust mass $\Sigma_{d}=1.18\times 10^{-26}\,\mathrm{g}\,N_{H}$ and adopting the extinction per hydrogen column density $A_{V}/N_{H}=3.24\times 10^{-22}$ mag cm2 from Table 2 of B. S. Hensley & B. T. Draine (2023), we find $\Sigma_{d}/A_{V}=0.2\,{\rm M}_{\odot}\,\mathrm{pc}^{-2}\,\mathrm{mag}^{-1}$. This estimate assumes that CGM dust has the same physical properties, e.g., composition and size distribution, as dust in the local ISM, an assumption which is unlikely to be accurate in detail (see discussion below). More refined constraints on the inferred dust mass would be possible if the full wavelength dependence of the extinction were to be determined in the future.
@@
There are some notable differences between the WISExSuperCOSMOS and LRG $A_{V}$ profiles, the most obvious of which is the sharp rise in the WISExSuperCOSMOS extinction at $r<50\ h^{-1}\,\mathrm{kpc}$. This excess is consistent with the inclusion in the WISExSuperCOSMOS sample of actively star-forming galaxies possessing dense inner CGM. Conversely, the LRG extinction profile has almost no dust at these close radii, which is consistent with LRGs having had no recent episodes of star formation. Although an inner $A_{V}$ excess was not evident in the original M10 result, numerous studies have since confirmed the broken power law shape of dust-induced absorption about star-forming (emission line) galaxies, inter alia, T.-W. Lan & H. Mo (2018); Z. Ruoyi & Y. Haibo (2020); Z. Chen et al. (2024). In fact, the WISExSuperCOSMOS profiles of Figure 3 bear a striking resemblance to Figure 4 of Z. Chen et al. (2024), in which the EW(Mg II) profiles of ELGs and LRGs and are superimposed and which features a very similar transition from high to low absorption at $\sim 60$ $h^{-1}\,$kpc.
@@
Having evaluated the inner halos, we now compare the profiles at larger distance scales. In addition to exhibiting a simpler profile shape, the LRG extinction profile is systematically higher than that of the WISExSuperCOSMOS sample and remains detectable out to 30 Mpc, with a slight dip in $A_{V}$ at $R\approx 2\,h^{-1}\,\mathrm{Mpc}$, near the boundary between the one-halo and two-halo regimes. Meanwhile, the WISExSuperCOSMOS $A_{V}$ tapers off at $R\gtrsim 1\,h^{-1}$ Mpc. The difference in $A_{V}$ alone is difficult to interpret, since the density of dust in the CGM depends on stellar mass, redshift, and star formation rate. Moreover, the redMaGiC LRG sample ($\bar{z}=0.35$) spans a larger cosmological volume than the WISExSCOS sample ($\bar{z}=0.14$), implying that the LRG sample’s dust column density should also be larger. Even so, a few qualitative statements can be made. LRGs tend to reside in denser environments where interactions are more frequent (M. Postman & M. J. Geller, 1984; I. K. Baldry et al., 2006; O. Curtis et al., 2024), offering more mechanisms for dust deposition and shielding (e.g., A. Boselli et al., 2022). By contrast, star-forming galaxies such as those in the WISExSCOS sample are on average lower-mass and less likely to reside in dense environments, thus dispersing their dust into a smaller volume within a relatively empty IGM where the dust is ultimately destroyed. Indeed, when comparing Figure 3 with the Mg II surface mass profiles in Figure 5 of G. Zhu et al. (2014), our $A_{V,\,\mathrm{stacked}}$ resembles a one-halo profile, while $A_{V,\,\mathrm{LRG}}$ appears to be the superposition of one-halo and two-halo components.
@@
The transition at 2 $h^{-1}\,$pc in both the LRG and WISExSCOS dust-extinction profiles is particularly interesting in the context of the “tired wind” model for galactic outflows driven by star formation (C. Lochhaas et al., 2018; R. Kannan et al., 2021). In this framework, galactic wind-driven bubbles—hot, fast outflows powered by supernova energy injection—expand and cool over Gyr timescales, ultimately “hanging” in the outer halo. While the C. Lochhaas et al. (2018) model tracks the evolution of a single wind bubble for an individual galaxy—a good description of the high $A_{V}$ in the inner halo of star-forming WISExSCOS galaxies—the observed LRG extinction profile can be interpreted as the collective effect of many such bubbles, launched at different epochs by different galaxies, persisting in the CGM over cosmic timescales. Based on the theoretical expectation that bubble sizes grow slowly to hundreds of kiloparsecs and that entrained dust can survive for Gyr (H. M. Richie et al., 2024), it is physically consistent for LRGs to host ancient, radially outflowing wind bubbles at 2–3 Mpc. This is also close to the spike in Mg II absorber redshift space distortions and truncation radius shown in Y. Zu (2021).

### v2

@@ III.2 Star catalogs for systematics testing
To create a background star catalog (for cross-correlation with the WISExSuperCOSMOS foreground), we query the DES Y3 GOLD photometric catalog using the star selection criteria in Sevilla-Noarbe et al. (2021). An additional cut of $18<$ mag_r_corr <21<21 eliminates saturated stars and possible galaxy confusion. The final catalog sizes are 824,345 stars in the “signal” catalog and 9,892,452 stars in the “random” catalog. Because these star catalogs contain the same magnitude measures as the joined redMaGiC-Y3 GOLD galaxy catalog, comparison of results between the two is very straightforward.
IV Methods
This section presents the analysis pipeline we put in place to apply the formalism of Section II and so measure circumgalactic dust extinction profiles. The major steps of the pipeline are: intake and processing of foreground and background catalogs including random catalogs; defining a particular dust extinction model; dividing of galaxies into redshift bins and de-meaning of galaxy colors; per-galaxy calculation of the value and uncertainty of $A_{\mathrm{V}}$; calculation of $\xi_{NA_{\mathrm{V}}}(\theta)$; and finally, the fitting of a polynomial to the resulting extinction profile. Details are provided in the subsections below. The code developed to carry out this analysis, dusthalos, is hosted on GitHub [^4] https://github.com/mcclearyj/dusthalos/ and will be made publicly available upon acceptance of this manuscript.
IV.1 Catalog handling
@@
The specific implementation of this workflow for the galaxy and star catalogs used in our analysis is mostly covered in Section III, with a few additional details below.
Masking. Although HEALPix survey masks are available for the DES Y3 and WISExSuperCOSMOS catalogs, they are not available for Gaia. To ensure an exact match in sky coverage of the foreground and background catalogs, we create our own using coordinate information in the catalog and the healpy Python package (Zonca et al., 2019).
@@
Photometry. The redMaGiC catalogs we use in our analysis include celestial coordinates and redshifts but not magnitudes. As described in Section III, to obtain colors for the real redMaGiC catalogs, we join member galaxies to their counterparts in the Y3 GOLD catalog on the COADD_OBJECT_ID parameter. Random catalog entries lack a corresponding entry in the Y3 GOLD catalog and so are randomly assigned photometry from the appropriate redshift slice in the photometry-augmented redMaGiC galaxy catalog. The color-matched random catalogs resulting from this procedure are shown in Figure 2, which shows good agreement with the real data.
@@
IV.2 Extinction calculation and correlation
To obtain values for $d\mathrm{M}/d\rho$, the model for dust-grain induced reddening referenced in Equations 35, we use the comprehensive dust extinction law published by Gordon et al. (2023), hereafter G23. The G23 model requires specification of the wavelengths at which to compute the extinction and a value for $R_{\mathrm{V}}$. For the former, we use the effective wavelengths of the DECam filter set available through the Virtual Observatory’s Filter Profile Service, rescaled as prescribed in Table 6 of Schlafly and Finkbeiner (2011): $\left\{g,r,i,z\right\}=\left\{4796.6,6382.6,7769.0,9108.2\right\}$. For the latter, in the absence of a compelling reason to do otherwise, we adopt the common default of $R_{\mathrm{V}}=3.1$. (Further consideration is given to the choice of $R_{\mathrm{V}}$ in Section VI.)
@@ IV.2 Extinction calculation and correlation
We compute $\xi_{NA_{\mathrm{V}}}(\theta)$ using the galaxy-kappa cross-correlation method NKCorrelation provided in the TreeCorr package [^5] https://github.com/rmjarvis/TreeCorr (Jarvis et al., 2004). Though intended to correlate galaxy counts with convergence (i.e., $\kappa$) the code is flexible enough to accommodate any scalar quantity such as $A_{\mathrm{V}}$. All correlations are taken in 13 impact parameter bins, from 0.5 to 200 arcminutes, or about 20 kiloparsecs to 10 megaparsecs in the average foreground galaxy rest frame. This range matches the physical scales probed in M10 and related works. Error bars on extinction profiles are based on TreeCorr’s sample covariance estimation method, which is the simplest method available in that package to account for the combination of Poisson counting statistics, statistical variance of the scalar quantity, and sample variance across the observation. Finally, to facilitate comparison across galaxy samples and the literature, we finally fit power laws of the form $A_{\mathrm{V}}(r)=Cr^{\alpha}$ to the extinction profiles using weighted least squares.
Our pipeline enables several tests for systematic error. Substitution of star catalogs for galaxy foregrounds and backgrounds offers a valuable null test for our pipeline, as main-sequence stars do not have extended halos of dust. Similarly, stars embedded in the Milky Way should not display dust-induced color gradients as a function of projected distance from galaxies outside the Milky Way. Additionally, inverting foreground and background samples, i.e., using a high-redshift catalog as a foreground and a low-redshift catalog as a background, should result in a non-detection of reddening. We deploy all three of these tests in the course of our analysis. Additional systematic errors that could affect our results are considered in Appendix B.
V Results
@@
$A^{\mathrm{M10}}_{\mathrm{V}}(r)=\left(4.4\pm 1.1\right)\times 10^{-3}\left(\frac{r}{100h^{-1}\,\mathrm{kpc}}\right)^{-0.86\pm 0.19}$ (9)
@@
V.1 Galaxy extinction profiles
@@
The extinction profiles of WISExSuperCOSMOS galaxies are shown in Figure 3: the result from cross-correlation with the $z>0.5$ high-density redMaGiC sample is plotted in dark blue, and the result from cross-correlation with the high-redshift redMaGiC sample is plotted in light blue. Both extinction profiles follow a broken power law, with a rapid decrease from $A_{\mathrm{V}}\sim 0.3$ to $A_{\mathrm{V}}\sim 0.02$ in the inner 50 $h^{-1}\,$kpc, then a slower decline to $A_{\mathrm{V}}\sim 0.02$ at 1 $h^{-1}\,$Mpc. At the transition to the 2-halo regime ($r\gtrsim 2$ $h^{-1}\,$Mpc), the extinction profiles rapidly diminish to below our detection threshold ($A_{\mathrm{V}}\sim 10^{-5}$).
@@
We fit a power law to the WISExSuperCOSMOS extinction profiles in the 1-halo regime (50 $h^{-1}\,$kpc <r<1<r<1 $h^{-1}\,$Mpc), where the lower limit is set to the point where the redMaGiC LRG and WISExSuperCOSMOS extinction profiles intersect, and the upper limit is set to the point where the extinction profiles taper off. Although the selection functions of the two redMaGiC background catalogs are different, the foreground catalog is of course the same. Accordingly, we fit the two WISExSuperCOSMOS $\times$ redMaGiC profiles with a single power law, which yields Equation 10. Quoted uncertainties are the standard error on the mean parameter values.
$A_{\mathrm{V,\,stacked}}(r)=\left(3.1\pm 0.06\right)\times 10^{-3}\left(\frac{r}{100\,h^{-1}\,{\rm kpc}}\right)^{-0.87\pm 0.04}$ (10)
@@
Cross-correlation of the redMaGiC LRG galaxy sample with the redMaGiC high-luminosity high-$z$ sample yields the extinction profile plotted in maroon in Figure 3. As might be expected for this galaxy population, there is little apparent dust at the lowest impact parameters, within a few optical radii of the galaxies ($r<50$ $h^{-1}\,$kpc). Between $0.05<r<2$ $h^{-1}\,$Mpc, the LRG profile is qualitatively similar to the WISExSCOS profiles. However, the LRG sample shows significant extinction out to 30 $h^{-1}\,$Mpc, with a distinct dip at $\sim 2$ $h^{-1}\,$Mpc—the point at which the WISExSCOS profiles taper off. We fit a power law to the LRG $A_{\mathrm{V}}$ profile in the range 50 $h^{-1}\,$kpc <r<30<r<30 $h^{-1}\,$Mpc, with the result given in Equation 11. As before, quoted uncertainties reflect the standard error on the mean parameter values.
$A_{\mathrm{V,\,LRG}}(r)=\left(4.6\pm 0.2\right)\times 10^{-3}\left(\frac{r}{100\,h^{-1}\,\mathrm{kpc}}\right)^{-0.77\pm 0.04}$ (11)
V.2 Systematics tests
@@
Figure 4 presents the results of the systematics tests described in Section IV, namely, the calculation of extinction profiles using star catalogs and with inverted foreground and background samples (high-$z$ foreground, low-$z$ background). The top row of Figure 4 shows the the extinction profile resulting from the cross-correlation of the Gaia star foreground catalog with the redMaGiC high-luminosity high-redshift sample (maroon points) and high-density sample (red points). The middle row of Figure 4 shows the $A_{\mathrm{V}}$ profile resulting from the cross-correlation of the WISExSuperCOSMOS foreground with the background catalog created from DES Y3 stars. For the reverse $A_{\mathrm{V}}$ profile test, we invert the redMaGiC LRG measurement and cross-correlate the high-redshift redMaGiC catalog with the $z<0.45$ high-density redMaGiC catalog. The outcome is shown in the bottom row of Figure 4.
@@
As with the galaxy samples, we fit power laws to the systematics-test extinction profiles, including the combined Gaia-redMaGiC high-density and high-redshift profiles. The resulting extinction profile parameters are summarized in Table 1 and also plotted in Figure 4. Weighted least squares fits to these extinction profiles are effectively all consistent with zero, with prefactors close to unity and exponents on the order of $10^{-5}$. The only fit that yielded a moderately significant ($3\,\sigma$) exponent was for the inverted foreground-background $A_{\mathrm{V}}$ profile. Even so, the magnitude of this exponent is only one ten-thousandth of the magnitude of the galaxy $A_{\mathrm{V}}$ profile exponents, indicating no cause for concern.
Table 1: Systematics test $A_{\mathrm{V}}$ profile fits
Sample Prefactor Exponent ($\times 10^{-5}$)
Gaia $\times$ redMaGiC high-density $\pm\,(8.0\times 10^{-5}) 1.0±-1.0\,\pm$ 1.7
Gaia $\times$ redMaGiC high-$z$ $1.0\,\pm\,(9.7\times 10^{-5})$ $0.3\,\pm\,2.0$
Gaia $\times$ redMaGiC combined $1.0\,\pm\,(6.0\times 10^{-5}$) $0.7\pm 1.3$
WISExSCOS $\times$ DES stars $1.0\,\pm\,(3.0\times 10^{-4}$) $5.0\,\pm\,8.3$
redMaGiC high-$z$ $\times$ low-$z$ $1.0\,\pm\,(1.3\times 10^{-4}$) $9.4\,\pm\,3.0$
@@ VI Discussion
Figure 3 shows an unambiguous detection of $A_{\mathrm{V}}$ at megaparsec scales about both redMaGiC LRGs and a mixed sample of quiescent and star-forming galaxies in WISExSCOS, with negligible contributions from systematic errors (Figure 4 and Table 1). The slopes ($\alpha\simeq{0.8}$) of the extinction profiles presented in Equations 10 and 11 are in excellent agreement with M10 (cf. Equation 9) and other studies both EW(Mg II)-based (Zhu et al., 2014; Lan and Mo, 2018; Chen et al., 2024) and extinction-based (Ruoyi and Haibo, 2020). This ubiquitous value of $\alpha\simeq 0.8$ is directly connected to halo mass profiles, where the so-called differential surface density typically follows $\Delta\Sigma\propto r^{\beta}$. For massive, dynamically relaxed galaxies, $0.7\leq\beta\leq 0.9$ (Brimioulle et al., 2013; Velander et al., 2014; van Uitert et al., 2016). These values of $\beta$ match both our measured $\alpha\simeq 0.8$ and values of $\alpha$ in the literature, which implies that dust permeates the entire galaxy halo. In addition to the exponent, the normalization of our $A_{\mathrm{V,\,LRG}}$ is also in excellent agreement with M10, while the $A_{\mathrm{V,\,stacked}}$ is about $1\,\sigma$ lower. Although a direct comparison of our result to Mg II absorbers is impossible without spectroscopic calibration of our samples, by applying the Ménard et al. (2008) scaling between E(BV)E(B-V) and EW(Mg II), we find a (very) approximate EW of 1.4 Å at an impact parameter of 20 kpc for our WISExSCOS sample, within a factor of 2 of the ELG results of Lan and Mo (2018) and Chen et al. (2024).
@@
We may convert our observed $A_{\mathrm{V}}$ profiles into an approximate dust mass surface density profile $\Sigma_{d}(r)$ using the astrodust+PAH model of Milky Way dust described in Hensley and Draine (2023). Using the astrodust surface dust mass $\Sigma_{d}=1.18\times 10^{-26}\,\mathrm{g}\,N_{H}$ and adopting the extinction per hydrogen column density $A_{\mathrm{V}}/N_{H}=3.24\times 10^{-22}$ mag cm2 from Table 2 of Hensley and Draine (2023), we find $\Sigma_{d}/A_{\mathrm{V}}=0.2\,{\rm M}_{\odot}\,\mathrm{pc}^{-2}\,\mathrm{mag}^{-1}$. This estimate assumes that CGM dust has the same physical properties, e.g., composition and size distribution, as dust in the local ISM, an assumption which is unlikely to be accurate in detail (see discussion below). More refined constraints on the inferred dust mass would be possible if the full wavelength dependence of the extinction were to be determined in the future.
@@
There are some notable differences between the WISExSuperCOSMOS and LRG $A_{\mathrm{V}}$ profiles, the most obvious of which is the sharp rise in the WISExSuperCOSMOS extinction at $r<50\ h^{-1}\,\mathrm{kpc}$. This excess is consistent with the inclusion of star-forming galaxies with dense inner CGM in the WISExSuperCOSMOS sample. Conversely, the LRG extinction profile has essentially no dust at these radii, which is consistent with LRGs having had no recent episodes of star formation. Although an inner $A_{\mathrm{V}}$ excess was not evident in the original M10 result, numerous studies have since confirmed the broken power law shape of dust-induced absorption about star-forming (emission line) galaxies, inter alia, Lan and Mo (2018); Ruoyi and Haibo (2020); Chen et al. (2024). In fact, the WISExSuperCOSMOS profiles of Figure 3 bear a striking resemblance to Figure 4 of Chen et al. (2024), in which the EW(Mg II) profiles of ELGs and LRGs and are superimposed and which features a very similar transition from high to low absorption at $\sim\!60$ $h^{-1}\,$kpc.
@@
At larger angular scales, the amplitude of the LRG extinction profile is systematically higher than the WISExSuperCOSMOS sample and remains detectable out to 30 Mpc, with a slight dip in $A_{\mathrm{V}}$ at $R\approx 2\,h^{-1}\,\mathrm{Mpc}$, near the boundary between the one-halo and two-halo regimes. Meanwhile, the WISExSuperCOSMOS $A_{\mathrm{V}}$ tapers off at $R\gtrsim 1\,h^{-1}$ Mpc. The difference in $A_{\mathrm{V}}$ alone is difficult to interpret, since the density of dust in the CGM depends on stellar mass, redshift, and star formation rate. Moreover, the redMaGiC LRG sample ($\bar{z}=0.35$) spans a larger cosmological volume than the WISExSCOS sample ($\bar{z}=0.14$), implying that the LRG sample’s dust column density should also be larger. Even so, a few qualitative statements can be made. LRGs tend to reside in denser environments where interactions are more frequent (Postman and Geller, 1984; Baldry et al., 2006; Curtis et al., 2024), offering more mechanisms for dust deposition and shielding (e.g., Boselli et al., 2022). By contrast, star-forming galaxies such as those in the WISExSCOS sample are on average lower-mass and less likely to reside in dense environments, thus dispersing their dust into a smaller volume within a relatively empty IGM where the dust is ultimately destroyed. Indeed, when comparing Figure 3 with the Mg II surface mass profiles in Figure 5 of Zhu et al. (2014), our $A_{V,\,\mathrm{stacked}}$ resembles a one-halo profile, while $A_{V,\,\mathrm{LRG}}$ appears to be the superposition of one-halo and two-halo components.
@@
The transition at 2 $h^{-1}\,$Mpc in both the LRG and WISExSCOS dust-extinction profiles is particularly interesting in the context of the “tired wind” model for galactic outflows driven by star formation (Lochhaas et al., 2018; Kannan et al., 2021). In this framework, galactic wind-driven bubbles—hot, fast outflows powered by supernova energy injection—expand and cool over Gyr timescales, ultimately “hanging” in the outer halo. While the Lochhaas et al. (2018) model tracks the evolution of a single wind bubble for an individual galaxy—a good description of the high $A_{\mathrm{V}}$ in the inner halo of star-forming WISExSCOS galaxies—the observed LRG extinction profile can be interpreted as the collective effect of many such bubbles, launched at different epochs by different galaxies, persisting in the CGM over cosmic timescales. Based on the theoretical expectation that bubble sizes grow slowly to hundreds of kiloparsecs and that entrained dust can survive for Gyr (Richie et al., 2024), it is physically consistent for LRGs to host ancient, radially outflowing wind bubbles at 2–3 Mpc. This is also close to the spike in Mg II absorber redshift space distortions and truncation radius shown in Zu (2021).